OpenStudy (anonymous):

what is the 100th term on the following sequence : 25 , 36 , 49 , 64 ,

OpenStudy (anonymous):

Do you see a pattern in the numbers of the sequence?

OpenStudy (anonymous):

Nope :lol

OpenStudy (anonymous):

Those four numbers don't look familiar at all?

OpenStudy (anonymous):

Mmmm..... nope *Thoguht Hard*

OpenStudy (anonymous):

I'll give you a hint. The first one is five times five.

OpenStudy (anonymous):

Oooooooo . Then 6*6 Ect ...

OpenStudy (anonymous):

Right. They're perfect squares.

OpenStudy (anonymous):

But the answeres Choices is 10,404 10,609 , 10,816 11,025

OpenStudy (anonymous):

You need the 100th term of the sequence. How could you describe the 100th term of the sequence?

OpenStudy (anonymous):

Mmm... x*x = 1000

OpenStudy (anonymous):

What? No. The first term of the sequence is 5*5, the second is 6*6, the third is 7*7, the hundredth is ...?

OpenStudy (anonymous):

\[ 5^2, 6^2,7^2, 8^2, \cdots 104^2 , \cdots \]

OpenStudy (anonymous):

104 ^2 = 10,816 and were did 104 come from ?

OpenStudy (anonymous):

\[ 104-5+1 =100 \]

OpenStudy (anonymous):

That's what I'm trying to get you to figure out. You need to look at the pattern. Your first term is 5^2, your second term is 6^2, your third term is 7^2, why is the hundredth term 104^2?

OpenStudy (anonymous):

6-5+1=2 is the second 7-5+1= 3 is the third 104-5+1=100 is the 100th

OpenStudy (anonymous):

\[ 1:5^2\\ 2:6^2\\ 3:7^2\\ 4:8^2\\ 100:104^2 \]

OpenStudy (anonymous):

The pattern for the \(n\)th term is \((n+4)^2\)

OpenStudy (anonymous):

Okay ill Study It *Snaped a Shoot of how to do it *

OpenStudy (anonymous):

.. Thanks